The decomposition of NH3 on platinum surface is zero order reaction. What are the rates of production of N2 and H2 if k = 2.5 × 10–4 mol–1 L s–1?

Grade:8

1 Answers

Gaurav
9 years ago
k = 2.5 × 10−4mol−1L s−1
Order of reaction = 0
Decomposition reaction of NH3
Here NH3 has 2 mol then N2 and H2 has 1 and 3 moles respectively
Use the formula of rate of reaction
Use this formula we get
Another formula of rate, Rate = k[NH3]x
Given that order of reaction x = 0
Plug the values we get
Plug the value of k we get
Rate of production of N2= k = 2.5 × 10−4mol L-1s−1
Rate of production of H2= 3k = 3 x 2.5 × 10−4mol L-1s−1= 7.5 × 10−4mol L-1s−1